OpenStudy (anonymous):

Given the function f(x) = The quantity of 3x minus 4, divided by 5, which of the below expressions is correct? f−1(x) = The quantity of 5x plus 4, divided by 3. f−1(x) = The quantity of 5x minus 4, divided by 3. f−1(x) = The quantity of negative 3x minus 4, divided by 5. f−1(x) = The quantity of 4 minus 3x, divided by 5. please can someone give me the answer.

OpenStudy (whpalmer4):

\[y = f(x) = \frac{3x-4}{5}\]To find \(f^{-1}(x)\) we swap the variables and solve for \(y\). \[x = \frac{3y-4}{5}\]\[y = f^{-1}(x) = \]

OpenStudy (whpalmer4):

Won't give you the answer, but I'll give you something much more useful: the knowledge of how to find the answer yourself.

OpenStudy (anonymous):

5x+4/3 is what I got. Is that right

OpenStudy (whpalmer4):

If you mean \[f^{-1}(x) = \frac{5x+4}{3}\]then that is correct, but as you wrote it, that is incorrect. You need to write it as \[(5x+4)/3\] if you are going to write it on one line.

OpenStudy (anonymous):

thanks